# Active Healing of Microtubule-Motor Networks

### v1

@@ Acknowledgements.
We are grateful to Zhen-Gang Wang and Howard A. Stone for fruitful discussions and to Inna Strazhnik for illustration of Fig. 4. This project is funded by NIH 1R35 GM118043 (MIRA), Packard Foundation, Moore Foundation, and Donna and Benjamin M. Rosen Bioengineering Center. FY also acknowledges financial support from BBE Divisional Postdoctoral Fellowship at Caltech.
@@ I.1 Overview
The main chemical components in our system are microtubules, motors and ATPs. The dynamics of active matter and solvent flows is initiated by a cascade of chemical reactions. Firstly, the monomer motors are activated by light and reversibly dimerized. The dimer motors then crosslink microtubules into a contracting network, which further drives a solvent flow. In our modeling, microtubules and motors are classified according to their chemical structures, such as monomer or dimer motors, and also to their crosslinking status, such as bound dimer motors and free dimer motors, as well as crosslinked and free microtubules. Here ”bound” means the motors are attached to the crosslinked microtubules, and ”free” means the microtubules or motors are not attached to the crosslinked network.
All the chemical reactions mentioned above, namely the reversible dimerization of motors and the crosslinking of microtubules can be modeled in continuity equations of these chemical components. The light patterns enter our model through the dimerization rate $p^{d}(x,y,t)$, where $x$ and $y$ are the spatial coordinates and $t$ is time. The shape of the light region can thereby be determined by a spatiotemporal function $p^{d}(x,y,t)$.
@@
Momentum equations determine the motion of active matter and solvent from their stresses and material properties. In principle, we can write down momentum equations for each chemical species. However, small proteins such as free (monomer and dimer) motors have little mechanical impact on the dynamics and can be assumed to follow the solvent flow. The crosslinked microtubules and bound dimer motors can be modeled together as a single “phase”, which is termed the “crosslinked network”. The crosslinked network is modeled as a viscoelastic gel, which self-contracts driven by its internal active stresses; the solvent flow is modeled as a Stokes flow driven by the contraction of the active gel and balanced by the hydrodynamic resistance in the flow cell; the free microtubules are modeled as passive particles. Each of the three phases exert friction on the other two phases. A summary of the three phases is shown in Fig. S1.
@@ I.2 Continuity Equations
Figure S2: Illustrations of chemical reactions in the model. A, Reversible dimerization of motors. B, Reversible crosslinking of a freely-moving microtubule $i$ and another microtubule $j$. The latter can be either free or pre-crosslinked. C, Reversible binding of a dimer motor on two pre-crosslinked microtubules.
The chemical reactions in our system are sketched in Fig. S2. The motor proteins can reversibly dimerize and the continuity equation is
@@
where $m$ is the monomer motor concentration, $d_{f}$ is the free dimer concentration, $\bm{u}$ is the solvent flow velocity, $D_{m}$ is the diffusivity, $p^{m}$ and $p^{d}(x,y,t)$ are the monomerization and dimerization rates, respectively. The function $p^{d}(x,y,t)$ is determined by the light pattern. In experiments, we use V-shaped light patterns, which can be defined by a set of points $A_{l}=\{(x,y)\in\mbox{light pattern}\}$. The light pattern is projected periodically onto the flow cell with time interval $T$ and duration $\Delta t$. Naively, the dimerization rate can be written as
@@
The continuity equations for the free ($d_{f}$) and the bound ($d_{b}$) dimer motors are
@@
where $d_{b}$ is the bound dimer concentration, $p_{2}^{\text{on}}$ is the rate of the dimer motors binding on the crosslinked microtubules, $D_{d}$ is the diffusivity of the free dimers, $p^{\text{off}}_{2}$ is the unbinding rate of a dimer motor that does not uncrosslink the two microtubules, see Fig. S2C. All free motors are assumed to follow the solvent flow ($\bm{u}$) and the bound motors follow the crosslinked gel ($\bm{v}$) .
@@
where $k_{A}$ is the consumption rate of the ATP by the bound dimer motors, and $D_{A}$ is the ATP diffusivity.
@@ I.3.1 Crosslinked network as an active gel
where $c$, $c_{A}$ and $d_{b}$ are the concentrations of the crosslinked microtubules, ATP and the bound dimer motors, respectively, $\alpha$ is a constant activity coefficient and $\bm{I}$ is the identity tensor. This active stress is derived to be $\bm{\sigma}^{\prime}_{a}\propto c^{2}\bm{I},$ in Ref [14], where the bound motor concentration is implicitly assumed to be uniform and does not appear in active stresses. In our experiments, the concentration difference of the dimer motors inside and outside of the illuminated regions is the key to form active gels. We can modify the theory in Ref [14] by adding a prefactor $d_{b}/c$ in the active stresses, which is the number of bound motors per microtubule.
@@ I.4 Non-dimensionalization
Using the initial microtubule concentration $c_{0}$, initial monomer motor concentration $m_{0}$, initial ATP concentration $c_{A0}$, and the typical length scale of the illuminated region $l$, we can non-dimensionalize the governing equations by
@@
The dimensionless continuity equations for the monomer (S4) and dimer (S9) motors are
@@ I.5 Simulations
Figure S3: Normalized concentration contours of crosslinked microtubules, free microtubules and free dimer motors. The contours of crosslinked microtubules (a) are encapsulated by the contours of of free microtubules (b) and free dimer motors (c) at the same levels, demonstrating that there is a boundary layer of free microtubules and motors outside the network surface. (e) and (f) are overlaid contours of crosslinked and free microtubules, and crosslinked and free dimer motors, respectively.
@@ I.5.1 Boundary layer of free microtubules and motors
We plot normalized concentrations of crosslinked microtubules, free microtubules and free dimer motors in Fig. S3. The contours are simulated results of Fig 1(b) in the main text at $\overline{t}=3$. From Fig. S3 we can see that the contours of crosslinked microtubules are always encapsulated by the contours of free microtubules and motors, confirming that there is a boundary layer at the surface of active networks.
@@ II.5 Microscope setup
We conducted the experiments using an automated widefield epifluorescence microscope (Nikon Ti-2), custom-modified for two additional imaging modes: epi-illuminated pattern projection and LED-gated transmitted light. Light patterns from a programmable DLP chip (EKB Technologies DLP LightCrafter™ E4500 MKII™ Fiber Couple) were projected onto the sample via a user-modified epi-illumination attachment (Nikon T-FL). The DLP chip was illuminated by a fiber-coupled 470 nm LED (ThorLabs M470L3). The epi-illumination attachment featured two light-path entry ports: one for the projected pattern light path and the other for a standard widefield epi-fluorescence light path. These light paths were combined using a dichroic mirror (Semrock BLP01-488R-25). The magnification of the epi-illumination system was calibrated to ensure the camera’s imaging sensor (FliR BFLY-U3-23S6M-C) was fully illuminated when the entire DLP chip was activated. Micro-Manager software, running custom scripts, controlled the pattern projection and stage movement. For the transmitted light path, we replaced the standard white-light brightfield source (Nikon T-DH) with an electronically time-gated 660 nm LED (ThorLabs M660L4-C5) to minimize light-induced dimerization during brightfield imaging.

### v2

@@ Acknowledgements.
We are grateful to Zhen-Gang Wang and Howard A. Stone for fruitful discussions and to Inna Strazhnik for illustration of Fig. 5. This project is funded by NIH 1R35 GM118043 (MIRA), Packard Foundation, Moore Foundation, and Donna and Benjamin M. Rosen Bioengineering Center. FY also acknowledges support from BBE Divisional Postdoctoral Fellowship at Caltech.
@@ I.1 Overview
The main chemical components in our system are microtubules, motors and ATPs. The dynamics of active matter and solvent flows is initiated by a cascade of chemical reactions. Firstly, the motors are activated by light and reversibly linked. The linked motors then crosslink microtubules into a contracting network, which further drives a solvent flow. In our modeling, microtubules and motors are classified according to their chemical structures, such as unlinked or linked motors, and also to their crosslinking status, such as bound linked motors and free linked motors, as well as crosslinked and free microtubules. Here ”bound” means the motors are attached to the crosslinked microtubules, and ”free” means the microtubules or motors are not attached to the crosslinked network.
All the chemical reactions mentioned above, namely the reversible “linking” of motors and the crosslinking of microtubules can be modeled in continuity equations of these chemical components. The light patterns enter our model through the dimerization rate $p^{d}(x,y,t)$, where $x$ and $y$ are the spatial coordinates and $t$ is time. The shape of the light region can thereby be determined by a spatiotemporal function $p^{d}(x,y,t)$.
@@
Momentum equations determine the motion of active matter and solvent from their stresses and material properties. In principle, we can write down momentum equations for each chemical species. However, small proteins such as free motors have little mechanical impact on the dynamics and can be assumed to follow the solvent flow. The crosslinked microtubules and bound linked motors can be modeled together as a single “phase”, which is termed the “crosslinked network”. The crosslinked network is modeled as a viscoelastic gel, which self-contracts driven by its internal active stresses; the solvent flow is modeled as a Stokes flow driven by the contraction of the active gel and balanced by the hydrodynamic resistance in the flow cell; the free microtubules are modeled as passive particles. Each of the three phases exert friction on the other two phases. A summary of the three phases is shown in Fig. S1.
@@ I.2 Continuity equations
Figure S2: Illustrations of chemical reactions in the model. A, Reversible linking of motors. B, Reversible crosslinking of a freely-moving microtubule $i$ and another microtubule $j$. The latter can be either free or pre-crosslinked. C, Reversible binding of a linked motor on two pre-crosslinked microtubules.
The chemical reactions in our system are sketched in Fig. S2. The motor proteins can reversibly link and the continuity equation is
@@
where $m$ is the unlinked motor concentration, $d_{f}$ is the free linked motor concentration, $\bm{u}$ is the solvent flow velocity, $D_{m}$ is the diffusivity, $p^{m}$ and $p^{d}(x,y,t)$ are the monomerization and dimerization rates, respectively. The function $p^{d}(x,y,t)$ is determined by the light pattern. In experiments, we use V-shaped light patterns, which can be defined by a set of points $A_{l}=\{(x,y)\in\mbox{light pattern}\}$. The light pattern is projected periodically onto the flow cell with time interval $T$ and duration $\Delta t$. Naively, the dimerization rate can be written as
@@
The continuity equations for the free ($d_{f}$) and the bound ($d_{b}$) linked motors are
@@
where $d_{b}$ is the bound linked motor concentration, $p_{2}^{\text{on}}$ is the rate of the linked motors binding on the crosslinked microtubules, $D_{d}$ is the diffusivity of the free linked motors, $p^{\text{off}}_{2}$ is the unbinding rate of a linked motor that does not uncrosslink the two microtubules, see Fig. S2C. All free motors are assumed to follow the solvent flow ($\bm{u}$) and the bound motors follow the crosslinked gel ($\bm{v}$) .
@@
where $k_{A}$ is the consumption rate of the ATP by the bound linked motors, and $D_{A}$ is the ATP diffusivity.
@@ I.3.1 Crosslinked network as an active gel
where $c$, $c_{A}$ and $d_{b}$ are the concentrations of the crosslinked microtubules, ATP and the bound linked motors, respectively, $\alpha$ is a constant activity coefficient and $\bm{I}$ is the identity tensor. This active stress is derived to be $\bm{\sigma}^{\prime}_{a}\propto c^{2}\bm{I},$ in Ref [19], where the bound motor concentration is implicitly assumed to be uniform and does not appear in active stresses. In our experiments, the concentration difference of the linked motors inside and outside of the illuminated regions is the key to form active gels. We can modify the theory in Ref [19] by adding a prefactor $d_{b}/c$ in the active stresses, which is the number of bound motors per microtubule.
@@ I.4 Non-dimensionalization
Using the initial microtubule concentration $c_{0}$, initial unlinked motor concentration $m_{0}$, initial ATP concentration $c_{A0}$, and the typical length scale of the illuminated region $l$, we can non-dimensionalize the governing equations by
@@
The dimensionless continuity equations for the unlinked (S4) and linked (S9) motors are
@@ I.5 Simulations
Figure S3: Normalized concentration contours of crosslinked microtubules, free microtubules and free linked motors. The contours of crosslinked microtubules (a) are encapsulated by the contours of of free microtubules (b) and free linked motors (c) at the same levels, demonstrating that there is a boundary layer of free microtubules and motors outside the network surface. (e) and (f) are overlaid contours of crosslinked and free microtubules, and crosslinked and free linked motors, respectively.
@@ I.5.1 Boundary layer of free microtubules and motors
We plot normalized concentrations of crosslinked microtubules, free microtubules and free linked motors in Fig. S3. The contours are simulated results of Fig 2(b) in the main text at $\overline{t}=3$. From Fig. S3 we can see that the contours of crosslinked microtubules are always encapsulated by the contours of free microtubules and motors, confirming that there is a boundary layer at the surface of active networks.
I.6 Force balance in the elastic-rod model.
Figure S4: (a) Schematics of the elastic-rod model, identical to Fig. 5 in the main text. (b) Schematics of the active healing force $F_{a}=\sigma dh/\sin\theta$ and the active compression $\sigma_{a}wd$ on the right arm. The viscous force exerted by the background solution on the arm is not plotted.
To illustrate how forces act on each arm in the elastic-rod model, we depict the active healing force $F_{h}=\sigma dh/\sin\theta$ and the active compression $\sigma_{a}wd$ in Fig. S4.
I.7 Motor activity and active healing
Figure S5: Simulations with different activities $\alpha$ show that the activity does not affect the arm curvatures but can change the time scales and densities of the active networks. Simulation in (a) is the same as Fig. 2(c) in the main text and used as a benchmark. Increasing (b) or decreasing (c) the activity $\alpha$ can significantly speed up or slow down the active matter dynamics, respectively. (d) At late times, smaller activity yields a loosely crosslinked network. (e) Depletion of ATP corresponding to (d). The ATP concentration is non-dimensionalized by the initial ATP concentration.
How does activity influence the healing process? The active stress $\bm{\sigma}_{a}$ in our simulations is modeled as an isotropic contractile stress $\sigma_{a}=\alpha c_{A}d_{b}c\bm{I}$, where $\alpha$ is the activity, $c_{A}$ is the ATP concentration, and $d_{b}$ is the concentration of motors bound to the active network. We numerically study how the activity $\alpha$ affects the bifurcation dynamics. Using the simulation in Fig. 2(c) in the main text as a benchmark, we find that decreasing activity can slow down the active matter dynamics (Fig. S5). The smaller activity also yields a more loosely crosslinked network at late times, as shown in Fig. S5(d), in contrast with the dense networks in Fig. 2(c) in the main text. However, the activity does not affect the arm curvatures, as seen from Fig. S5(a-c). We also vary the activity on networks with different opening angles and get the same conclusions. The unvarying curvatures indicate that changing activity does not impact the competition of the merging and buckling forces and therefore does not change the critical angle $\theta^{*}$ in bifurcation. This is not surprising as both the merging and compression forces originate from the same microscopic active stress $\bm{\sigma}_{a}$. Our simulations suggest that increasing activity can enhance both healing and compression forces in a similar manner such that even though the overall dynamics is accelerated, the critical angle $\theta^{*}$, which depends on the ratio of the compression and merging forces, remains largely unaffected. The loosely crosslinked network in Fig. S5(d) is a natural consequence of weaker contractile stresses. Additionally, the energy consumption, or equivalently, the ATP consumption also plays an important role in determining the late-stage network densities. In our simulations, the ATP consumption rate is modeled as $-k_{A}d_{b}c_{A}$ with a constant coefficient $k_{A}$ (see equation (S10)). Note that we assume $k_{A}$ is unrelated to the activity $\alpha$. Since a smaller activity prolongs the dynamics and consumes more ATP over a longer time span, at late times the depletion of ATP, as shown in Fig. S5(e), will stall further contraction of the network, which also gives rise to a more loosely crosslinked network.
@@ II.5 Microscope setup
We conducted the experiments using an automated widefield epifluorescence microscope (Nikon Ti-2), custom-modified for two additional imaging modes: epi-illuminated pattern projection and LED-gated transmitted light. Light patterns from a programmable DLP chip (EKB Technologies DLP LightCrafter™ E4500 MKII™ Fiber Couple) were projected onto the sample via a user-modified epi-illumination attachment (Nikon T-FL). The DLP chip was illuminated by a fiber-coupled 470 nm LED (ThorLabs M470L3). The epi-illumination attachment featured two light-path entry ports: one for the projected pattern light path and the other for a standard widefield epi-fluorescence light path. These light paths were combined using a dichroic mirror (Semrock BLP01-488R-25). The magnification of the epi-illumination system was calibrated to ensure the camera’s imaging sensor (FliR BFLY-U3-23S6M-C) was fully illuminated when the entire DLP chip was activated. Micro-Manager software, running custom scripts, controlled the pattern projection and stage movement. For the transmitted light path, we replaced the standard white-light brightfield source (Nikon T-DH) with an electronically time-gated 660 nm LED (ThorLabs M660L4-C5) to minimize light-induced linkage during brightfield imaging.
@@ II.6 Light pattern construction
Figure S6: Experimental images of concave networks with (a) $\theta=20^{\circ}$ and (b) $\theta=30^{\circ}$. Scale bars, 100 $\mu$m.
@@ II.7 Image processing for centerline-curvature measurement
II.8 Experimental images of concave networks
We show more examples of concave V-shaped networks with small angles in Fig. S6.